OpenStudy (anonymous):

Help!! WILL MEDAL~!! I HAVE TO FINISH THIS CLASS What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what are we using quadratic formula?

OpenStudy (anonymous):

first do inverse and do 4x2+3=-12x /4x /4x

OpenStudy (anonymous):

@dan815 help!!

OpenStudy (anonymous):

And this is quadratec functions... I don't understand them at all

OpenStudy (anonymous):

I know the answer

OpenStudy (anonymous):

u get -3x so 2+3=-3x

OpenStudy (anonymous):

@soul_reaper ok then what is it....

OpenStudy (anonymous):

5=-3x /-3 /-3 1 and 2/3=x

OpenStudy (anonymous):

0r 1.666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666

OpenStudy (anonymous):

im not here to give answer im here to help

OpenStudy (anonymous):

it also says numerical answers expected

OpenStudy (anonymous):

she needs it im guessing soon

OpenStudy (anonymous):

extremely soon.

OpenStudy (anonymous):

how soon?

OpenStudy (anonymous):

what is 1.6666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666666 to the nearest hundreth

OpenStudy (anonymous):

so there is no C?

OpenStudy (anonymous):

and like now if possible. I need to be done with this class by monday ....

OpenStudy (anonymous):

sorry trying to help two people at a time

OpenStudy (anonymous):

1.66 ?

OpenStudy (anonymous):

one time*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is that the answer??

OpenStudy (anonymous):

@DaBest

OpenStudy (anonymous):

@soul_reaper